OpenStudy (anonymous):

What is the value for x in the proportion x+8/5x-2=3/8 needdd help

OpenStudy (anonymous):

help please i been like 10 mins on this one problem !

OpenStudy (anonymous):

(x+8)/(5x-2)=3/8 8(x+8)=3(5x-2) 8x+64=15x-6 -7x=-70 x=10

OpenStudy (anonymous):

i see where i went wrong thanks

OpenStudy (anonymous):

No problem. :]